# Stochastic weight matrix dynamics during learning and Dyson Brownian motion

### v1

@@ 1 Introduction
Recent years have seen a dramatic increase of the use of machine learning (ML) in the fundamental sciences, with the adoption and development of many ML applications to improve and speed-up scientific analysis [1]. Reversely there is a growing trend to use the methodology of (theoretical) physics to understand ML algorithms, viewing these as acting on systems with many fluctuating degrees of freedom and hence employing the analogy with statistical physics. In this paper we explore this second direction and argue that learning can be formulated in the framework of Dyson Brownian motion [2], thereby inheriting many features of random matrix theory (RMT) [3, 4, 5, 6, 7, 8].
@@
In general terms, ML algorithms aim to minimise some cost function by applying stochastic gradient descent (SGD), or variations thereof, to weight matrices $W$ defined inside the architecture. Stochastic updates of matrices immediately establishes the link with Dyson Brownian motion, which exactly describes those, and yields an equation for the dynamics of the eigenvalues of the symmetric combination $X=W^{T}W$. This dynamics contains universal aspects, e.g. eigenvalue repulsion due to an induced Coulomb term, as reflected in the Wigner surmise, as well as non-universal aspects, related to details of the gradient of the loss function. In many algorithms, stochasticity is inherently present due sampling and mini-batch updates. We demonstrate that the distribution of eigenvalues depends on the ratio of the learning rate $\alpha$ and mini-batch size $|{\cal B}|$, and not on these quantities separately. This linear scale rule has been observed previously [9, 10, 11, 12], but we demonstrate that it is a direct consequence of stochastic matrix dynamics and Dyson Brownian motion. An interesting corollary is that there is no simple limit (e.g.  $\alpha\to 0$, $1/|{\cal B}|\to 0$) in which SGD reduces to a stochastic differential equation (SDE) in continuous time [13, 14]. Again, this has been noted before [15], but we emphasise the viewpoint that a tunable ratio $\alpha/|{\cal B}|$ in fact prevents an obvious SDE limit (in the limit $\alpha/|{\cal B}|\to 0$, the SGD update reduces to an ordinary differential equation for the eigenvalues). The amount of stochasticity is directly proportional to $\alpha/|{\cal B}|$ and therefore determines the strength of the eigenvalue repulsion. It hence sets a fundamental limit on the accuracy of learning, but reversely a larger level of stochasticity leads to better generalisation, as it avoids overfitting [9, 10, 11]. These concepts find a natural place in the framework of Dyson Brownian motion, as we will demonstrate.
@@
This paper is organised as follows. After setting up the general framework in Sec. 2, we demonstrate our arguments explicitly in two cases, namely the Gaussian restricted Boltzmann machine (RBM) and a simple teacher-student model in Sec. 3. The conclusions are summarised in Sec. 4, along with directions for future research. App. A contains a brief summary of Dyson Brownian motion, while Apps.  B and C contain some further comments used in the main text.
@@
With relation to RMT, we note that Refs. [16, 17] apply RMT to the distribution of eigenvalues of weight matrices in an empirical manner, but not in the framework of Dyson Brownian motion and the linear scaling rule. In Ref. [18] it was shown that “real-world data” have many features consistent with RMT. Here we apply RMT to weight matrices, not to data.
2 Stochastic weight matrix dynamics
2.1 Stochastic updates
@@
$W\to W^{\prime}=W+\delta W=W-\alpha\frac{\delta{\cal L}}{\delta W},$ (2.1)
@@
$\delta W_{\cal B}=\frac{1}{|{\cal B}|}\sum_{b\in{\cal B}}\delta W_{b}.$ (2.2)
The entire data set from which the mini-batches are drawn is assumed to be sufficiently large to not lead to additional constraints. Since the mini-batch size is finite, $\delta W_{\cal B}$ is a stochastic variable, with the size of the fluctuations set by the central limit theorem. We hence write the update as
$\delta W=\delta W_{\cal B}+\frac{1}{\sqrt{|{\cal B}|}}\sqrt{{\rm Var}(\delta W)}\,\eta,$ (2.3)
where the first term is the “deterministic” part of the update for a given mini-batch and the second part reflects stochastic fluctuations, whose magnitude decreases with increasing mini-batch size. The variance is defined for each matrix element, i.e.,
${\rm Var}(\delta W_{ij})=\langle\delta W_{ij}\delta W_{ij}\rangle-\langle\delta W_{ij}\rangle\langle\delta W_{ij}\rangle$ (2.4)
(no summation over repeated indices) and the noise $\eta_{ij}\sim{\cal N}(0,1)$ is matrix-valued. In terms of the gradient of the loss function, this update reads
$W^{\prime}=W-\alpha\left(\frac{\delta\cal L}{\delta W}\right)_{\cal B}+\frac{\alpha}{\sqrt{|{\cal B}|}}\sqrt{\mbox{Var}\left(\frac{\delta\cal L}{\delta W}\right)}\eta,$ (2.5)
@@
$X\to X^{\prime}=X+\delta W^{T}W+W^{T}\delta W\equiv X+\delta X,$ (2.7)
@@
$X\to X^{\prime}=X+\delta X_{\cal B}+\frac{1}{\sqrt{|{\cal B|}}}\sqrt{\mbox{Var}(\delta X)}\eta,$ (2.8)
as above.
2.2 Dyson Brownian motion
The framework in which to consider stochastic matrix dynamics for a symmetric matrix $X$ is Dyson Brownian motion [2], which is summarised in App. A. The main feature is that the eigenvalues of $X$ not only evolve stochastically but also repel, due to an induced Coulomb term. The strength of the stochastic term and of the Coulomb term are related to the stochasticity in the original matrix equation for $X$.
Starting from Eq. (2.8), the eigenvalues $x_{i}$ of $X$ evolve according to – see Eq. (A.2) –
$x_{i}\to x_{i}^{\prime}=x_{i}+K_{i}+\sum_{j\neq i}\frac{g_{i}^{2}}{x_{i}-x_{j}}+\sqrt{2}g_{i}\eta_{i},$ (2.9)
where $K_{i}$ and $g_{i}$ are linked to the deterministic and stochastic terms in Eq. (2.8), and again $\eta\sim{\cal N}(0,1)$. The term with the summation is the Coulomb term, resulting in eigenvalue repulsion. The learning rate and batch size can be made explicit by writing, c.f. Eq. (2.5),
$K_{i}=\alpha\tilde{K}_{i},\qquad g_{i}=\frac{\alpha}{\sqrt{|{\cal B|}}}\tilde{g}_{i},$ (2.10)
@@ 2.2 Dyson Brownian motion
$x_{i}\to x_{i}^{\prime}=x_{i}+\alpha\tilde{K}_{i}+\frac{\alpha^{2}}{|{\cal B}|}\sum_{j\neq i}\frac{\tilde{g}_{i}^{2}}{x_{i}-x_{j}}+\frac{\alpha}{\sqrt{|{\cal B}|}}\sqrt{2}\tilde{g}_{i}\eta_{i}.$ (2.11)
When trying to identify the learning rate or the inverse batch size with a stepsize $\epsilon$, we note here that the drift, including the Coulomb term, and the stochastic term do not scale in the standard (i.e. $\rm\hat{I}$to calculus) manner. Indeed, the difficulties of going from stochastic gradient descent to a stochastic differential equation (SDE) [13, 14] are well known [15], with the update reducing to an ordinary differential equation in the limit that e.g. the learning rate goes to zero. However, as we will demonstrate shortly, the appearance of the learning rate and batch size as in Eq. (2.11) naturally leads to the linear scaling rule, in which $\alpha/|{\cal B}|$ is a parameter whose tunability can be exploited. An obvious SDE limit should therefore in fact not be expected.
We will derive the linear scaling rule [10, 11] now, by considering the stationary distribution corresponding to the stochastic process (2.9). The associated Fokker-Planck equation for the distribution is given in App. A. The stationary distribution is known as the Coulomb gas, which reads
$P_{s}(\{x_{i}\})=\frac{1}{Z}\prod_{i<j}\left|x_{i}-x_{j}\right|e^{-\sum_{i}V_{i}(x_{i})/g_{i}^{2}},$ (2.12)
@@
Here it is assumed that the drift can be derived from a potential,
$K_{i}(x_{i})=-\frac{dV_{i}(x_{i})}{dx_{i}}.$ (2.14)
@@
$\frac{V_{i}(x_{i})}{g_{i}^{2}}=\frac{1}{\alpha/|{\cal B}|}\frac{\tilde{V}_{i}(x_{i})}{\tilde{g}_{i}^{2}},$ (2.16)
@@
$\tilde{V}_{i}(x_{i})=\tilde{V}_{i}(x_{i}^{s})+\frac{1}{2}\Omega_{i}\left(x_{i}-x_{i}^{s}\right)^{2}+\ldots,$ (2.17)
with $\Omega_{i}$ the curvature around the minimum, the exponentials in Eq. (2.12) are Gaussians centred at $x_{i}=x_{i}^{s}$ with variance
$\sigma_{i}^{2}=\frac{\alpha}{|{\cal B}|}\frac{\tilde{g}_{i}^{2}}{\Omega_{i}}.$ (2.18)
The interplay between learning rate, batch size, curvature of the loss function, and the variance of fluctuations is then easy to see. We emphasise that the assumption of a single well-defined minimum is not required in the derivation leading to Eq. (2.16).
We conclude that the stationary distribution after training depends on the ratio of the learning rate and batch size, and not on these quantities separately. This was observed empirically previously [10, 11], but we have demonstrated that it is a direct consequence of stochastic matrix dynamics, when cast in this framework.
The Coulomb gas structure of the stationary distribution has further, universal, consequences. Due to the Coulomb term, eigenvalues repel and the spacing between eigenvalues is nonzero. This effect is captured in the Wigner surmise, which is the distribution for the level spacings $S_{i}=x_{i+1}-x_{i}$, see Eqs. (A.11, A.13) . The spectral density,
$\rho(x)=\left\langle\frac{1}{N}\sum_{i=1}^{N}\delta(x-x_{i})\right\rangle,$ (2.19)
takes a universal form as well, called the Wigner semicircle, see Eq. (A.15). In the following we explore both universal and non-universal features in two examples.
3 Applications
@@ 3.1 Gaussian RBMs
Restricted Boltzmann machines [19, 20] consist of one visible layer (with $N_{v}$ nodes) and one hidden layer (with $N_{h}$ nodes), see Fig. 1. There are no connections within each layer. The degrees of freedom can be discrete, as in an Ising model, continuous or mixed; Ref. [21] is a useful review. We consider Gaussian RBMs, in which both sets of degrees of freedom are quadratic. The ones on the visible layer are collected in an $N_{v}$-dimensional vector $\phi$ and on the hidden layer in an $N_{h}$-dimensional vector $h$. They are coupled bilocally via the $N_{v}\times N_{h}$ weight matrix $W$. We follow the notation of our previous work [22].
@@
$E(\phi,h)=\frac{1}{2}\mu^{2}\phi^{T}\phi+\frac{1}{2\sigma_{h}^{2}}(h-\eta)^{T}(h-\eta)-\phi^{T}Wh.$ (3.2)
@@
$p(\phi)=\int Dh\,p(\phi,h)=\frac{1}{Z}\exp\left(-\frac{1}{2}\phi^{T}K\phi+J^{T}\phi\right),$ (3.3)
@@
$K=\mu^{2}1\!\!1-\sigma_{h}^{2}WW^{T},\qquad\qquad J=W\eta.$ (3.4)
Here $1/\mu^{2}$ is the variance on the visible layer, $\sigma_{h}^{2}$ is the variance on the hidden layer, and $\eta$ is a bias, which we will put to zero from now on. This model has been studied in detail from the perspective of lattice field theory (LFT) in Ref. [22], which explains the appearance of $\mu^{2}$ as a mass parameter for a scalar field. Its generative power is linked explicitly to the number of hidden nodes $N_{h}$ and the mass parameter $\mu^{2}$, which both act as an ultraviolet regulator, using the terminology familiar from LFT. To match to the general notation of the previous section, we identify $N_{v}=M$ and $N_{h}=N$ and take $N\leq M$.
@@
$\frac{\delta{\cal L}}{\delta W_{ia}}=\sigma_{h}^{2}\big{(}\langle\phi_{i}\phi_{j}\rangle_{\rm target}-\langle\phi_{i}\phi_{j}\rangle_{\rm model}\big{)}W_{ja}.$ (3.5)
Here the first two-point function is evaluated using the target data and the second one is the RBM prediction. The dynamics can be analysed semi-analytically [21, 22] by performing a singular-value decomposition $W=U\Xi V^{T}$ in which $U$ and $V$ are orthogonal transformations in $M$ and $N$ dimensions respectively, and $\Xi$ is a rectangular matrix with $N$ singular values $\xi_{i}$ ($i=1,\ldots,N$) on the diagonal.
@@
$D_{K}=\mbox{diag}\big{(}\underbrace{\mu^{2}-\sigma_{h}^{2}\xi_{1}^{2},\mu^{2}-\sigma_{h}^{2}\xi_{2}^{2},\ldots,\mu^{2}-\sigma_{h}^{2}\xi_{N}^{2}}_{N},\underbrace{\mu^{2},\ldots,\mu^{2}}_{M-N}\big{)}.$ (3.7)
@@
$\kappa_{k}=m^{2}+p_{{\rm lat},k}^{2}=m^{2}+2-2\cos\left(\frac{2\pi k}{N_{v}}\right),$ (3.8)
@@
The numerical training of the RBM is carried out using persistent contrastive divergence (PCD) with mini-batches, see Ref. [22] for details. The RBM mass parameter is fixed at $\mu^{2}=9$ and $N_{v}=N_{h}=10$. The learning rate and the mini-batch size can be independently varied. For each choice of learning rate and mini-batch size, we have trained 7500 RBMs, to gather statistics. A typical example of the learnt distribution of eigenvalues $x_{i}$ of $X$ is shown in Fig. 3.
@@
3.2 Universal predictions
The task is to describe the distributions shown in Fig. 3 using the concepts from random matrix theory presented above. We start with the universal predictions. Due to the eigenvalue repulsion, degenerate modes cannot be learnt exactly. This can be analysed using the Wigner surmise: for each of the four doubly-degenerate modes we collect data on the level splitting for adjacent eigenvalues, $S_{i}=x_{i+1}-x_{i}$. Here we employed the fact that the non-degenerate eigenvalues are well-separated. The resulting histograms are shown in Fig. 4 (left). These histograms can be fitted with a single-parameter Ansatz – see Eq. (A.11) – in terms of the mean level splitting $\langle S_{i}\rangle=\sqrt{\pi}\sigma_{i}$. Presenting the same data as a function of $s=S_{i}/\langle S_{i}\rangle$ yields the parameter-free Wigner surmise – see Eq. (A.13) – leading to a universal description of the level spacing. The linear rise at small $s$ reflects the level repulsion.
@@ 3.2 Universal predictions
Next we turn to the spectral density (A.14) for each of the four doubly-degenerate modes. The results for two of these are shown in Fig. 5 (note that these are the second-to-lowest and second-to-highest histograms previously shown in Fig. 3). The lines are fits to Eq. (A.15) with the position $x_{m,i}$ and width $\sigma_{i}$ as free parameters. It is clear that the histograms are not described by simple Gaussians, but are broader and flatter. Non-Gaussianity can be checked by computing the Binder cumulant $U_{4}$ of the distribution, which equals (here $\delta x=x-x_{m}$)
$U_{4}\equiv\frac{\left\langle\delta x^{4}\right\rangle}{3\left\langle\delta x^{2}\right\rangle^{2}}-1=-\frac{4}{27}\approx-0.148$ (3.9)
for the Wigner semi-circle (with $N=2$), while it vanishes for a Gaussian distribution.

### v2

@@ I Introduction
Recent years have seen a dramatic increase of the use of machine learning (ML) in the fundamental sciences, with the adoption and development of many ML applications to improve and speed-up scientific analysis Carleo_2019 . Reversely there is a growing trend to use the methodology of (theoretical) physics to understand ML algorithms, viewing these as acting on systems with many fluctuating degrees of freedom and hence employing the analogy with statistical physics. In this paper we explore this second direction and argue that learning can be formulated in the framework of Dyson Brownian motion Dyson-4 , thereby inheriting many features of random matrix theory (RMT) Wigner-1 ; Wigner-2 ; Dyson-1 ; Dyson-2 ; Dyson-3 ; Meh2004 .
@@
In general terms, ML algorithms aim to minimise some cost function by applying stochastic gradient descent (SGD), or variations thereof, to weight matrices $W$ defined inside the architecture. Stochastic updates of matrices immediately establishes the link with Dyson Brownian motion, which exactly describes those, and yields an equation for the dynamics of the eigenvalues of the symmetric combination $X=W^{T}W$. This dynamics contains universal aspects, e.g. eigenvalue repulsion due to an induced Coulomb term, as reflected in the Wigner surmise, as well as non-universal aspects, related to details of the gradient of the loss function. In many algorithms, stochasticity is inherently present due sampling and mini-batch updates. We demonstrate that the distribution of eigenvalues depends on the ratio of the learning rate $\alpha$ and mini-batch size $|{\cal B}|$, and not on these quantities separately. This linear scaling rule has been observed previously Goyal-1 ; Smith-1 ; Smith-2 ; Smith-3 (see also Ref. DBLP:journals/corr/abs-1710-11029 ), but we demonstrate that it is a direct consequence of stochastic matrix dynamics and Dyson Brownian motion. An interesting corollary is that there is no simple limit (e.g.  $\alpha\to 0$, $1/|{\cal B}|\to 0$) in which SGD reduces to a stochastic differential equation (SDE) in continuous time mandt2015 ; li2017 . Again, this has been noted before yaida2018 , but we emphasise the viewpoint that a tunable ratio $\alpha/|{\cal B}|$ in fact prevents an obvious SDE limit (in the limit $\alpha/|{\cal B}|\to 0$, the SGD update reduces to an ordinary differential equation for the eigenvalues). The amount of stochasticity is directly proportional to $\alpha/|{\cal B}|$ and therefore determines the strength of the eigenvalue repulsion. It hence sets a fundamental limit on the accuracy of learning, but reversely a larger level of stochasticity leads to better generalisation, as it avoids overfitting Goyal-1 ; Smith-1 ; Smith-2 . These concepts find a natural place in the framework of Dyson Brownian motion, as we will demonstrate.
@@
This paper is organised as follows. After setting up the general framework in Sec. II, we demonstrate our arguments explicitly in two cases, namely the Gaussian restricted Boltzmann machine (RBM) and a simple teacher-student model in Sec. III. The conclusions are summarised in Sec. IV, along with directions for future research. App. A contains a brief summary of Dyson Brownian motion, while Apps.  B, C and D contain some further comments used in the main text.
@@
With relation to RMT, we note that Refs. Martin-2019 ; Baskerville apply RMT to the distribution of eigenvalues of weight matrices in an empirical manner, but not in the framework of Dyson Brownian motion and the linear scaling rule. In Ref. Levi:2023qwg it was shown that “real-world data” have many features consistent with RMT. Here we apply RMT to weight matrices, not to data.
II Stochastic weight matrix dynamics
II.1 Stochastic updates
@@
$W_{ij}\to W_{ij}^{\prime}=W_{ij}+\delta W_{ij}=W_{ij}-\alpha\frac{\delta{\cal L}}{\delta W_{ij}},$ (1)
@@
$\delta W_{ij}^{\cal B}=\frac{1}{|{\cal B}|}\sum_{b\in{\cal B}}\delta W_{ij}^{b}.$ (2)
The entire data set from which the mini-batches are drawn is assumed to be sufficiently large to not lead to additional constraints. Since the mini-batch size is finite, $\delta W^{\cal B}$ is a stochastic variable, with the size of the fluctuations set by the central limit theorem. We hence write the update as
$\delta W_{ij}=\delta W_{ij}^{\cal B}+\frac{1}{\sqrt{|{\cal B}|}}\sqrt{{\rm Var}(\delta W_{ij})}\,\eta_{ij},$ (3)
where the first term is the “deterministic” part of the update for a given mini-batch and the second part reflects stochastic fluctuations, whose magnitude decreases with increasing mini-batch size. The noise $\eta_{ij}\sim{\cal N}(0,1)$ is matrix-valued. The variance is defined for each matrix element, i.e.,
${\rm Var}(\delta W_{ij})=\langle\delta W_{ij}\delta W_{ij}\rangle-\langle\delta W_{ij}\rangle\langle\delta W_{ij}\rangle.$ (4)
The update (3) is applied to each matrix element and no summation over repeated indices is implied in this section. In terms of the gradient of the loss function, this update reads
$W_{ij}^{\prime}=W_{ij}-\alpha\left(\frac{\delta\cal L}{\delta W_{ij}}\right)_{\cal B}+\frac{\alpha}{\sqrt{|{\cal B}|}}\sqrt{\mbox{Var}\left(\frac{\delta\cal L}{\delta W_{ij}}\right)}\eta_{ij},$ (5)
@@
$X\to X^{\prime}=X+\delta W^{T}W+W^{T}\delta W\equiv X+\delta X,$ (7)
@@
$X_{ij}\to X_{ij}^{\prime}=X_{ij}+\delta X_{ij}^{\cal B}+\frac{1}{\sqrt{|{\cal B|}}}\sqrt{\mbox{Var}(\delta X_{ij})}\eta_{ij},$ (8)
as above, with symmetric noise in this case.
II.2 Dyson Brownian motion
The framework in which to consider stochastic matrix dynamics for a symmetric matrix $X$ is Dyson Brownian motion Dyson-4 , which is summarised in App. A. The main feature is that the eigenvalues of $X$ not only evolve stochastically but also repel, due to an induced Coulomb term. The strength of the stochastic term and of the Coulomb term are related to the stochasticity in the original matrix equation for $X$.
Starting from Eq. (8), the eigenvalues $x_{i}$ of $X$ evolve according to – see Eq. (63) –
$x_{i}\to x_{i}^{\prime}=x_{i}+K_{i}+\sum_{j\neq i}\frac{g_{i}^{2}}{x_{i}-x_{j}}+\sqrt{2}g_{i}\eta_{i},$ (9)
where $K_{i}$ and $g_{i}$ are linked to the deterministic and stochastic terms in Eq. (8), and again $\eta\sim{\cal N}(0,1)$. The term with the summation is the Coulomb term, resulting in eigenvalue repulsion. The learning rate and batch size can be made explicit by writing, c.f. Eq. (5),
$K_{i}=\alpha\tilde{K}_{i},\qquad\qquad g_{i}=\frac{\alpha}{\sqrt{|{\cal B|}}}\tilde{g}_{i},$ (10)
@@ II.2 Dyson Brownian motion
$x_{i}\to x_{i}^{\prime}=x_{i}+\alpha\tilde{K}_{i}+\frac{\alpha^{2}}{|{\cal B}|}\sum_{j\neq i}\frac{\tilde{g}_{i}^{2}}{x_{i}-x_{j}}+\frac{\alpha}{\sqrt{|{\cal B}|}}\sqrt{2}\tilde{g}_{i}\eta_{i}.$ (11)
When trying to identify the learning rate or the inverse batch size with a stepsize $\epsilon$, we note here that the drift, including the Coulomb term, and the stochastic term do not scale in the standard (i.e. Itô calculus) manner. Indeed, the difficulties of going from stochastic gradient descent to a stochastic differential equation (SDE) mandt2015 ; li2017 are well known yaida2018 , with the update reducing to an ordinary differential equation in the limit that e.g. the learning rate goes to zero. However, as we will demonstrate shortly, the appearance of the learning rate and batch size as in Eq. (11) naturally leads to the linear scaling rule, in which $\alpha/|{\cal B}|$ is a parameter whose tunability can be exploited. An obvious SDE limit should therefore in fact not be expected.
We will derive the linear scaling rule Smith-1 ; Smith-2 now, by considering the stationary distribution corresponding to the stochastic process (9). The associated Fokker-Planck equation for the distribution is given in App. A. The stationary distribution is known as the Coulomb gas, which reads
$P_{s}(\{x_{i}\})=\frac{1}{Z}\prod_{i<j}\left|x_{i}-x_{j}\right|e^{-\sum_{i}V_{i}(x_{i})/g_{i}^{2}},$ (12)
@@
Here it is assumed that the drift can be derived from a separable potential,
$K_{i}(x_{i})=-\frac{dV_{i}(x_{i})}{dx_{i}}.$ (14)
@@
$\frac{V_{i}(x_{i})}{g_{i}^{2}}=\frac{1}{\alpha/|{\cal B}|}\frac{\tilde{V}_{i}(x_{i})}{\tilde{g}_{i}^{2}},$ (16)
@@
$\tilde{V}_{i}(x_{i})=\tilde{V}_{i}(x_{i}^{s})+\frac{1}{2}\Omega_{i}\left(x_{i}-x_{i}^{s}\right)^{2}+\ldots,$ (17)
with $\Omega_{i}$ the curvature around the minimum, the exponentials in Eq. (12) are Gaussians centred at $x_{i}=x_{i}^{s}$ with variance
$\sigma_{i}^{2}=\frac{\alpha}{|{\cal B}|}\frac{\tilde{g}_{i}^{2}}{\Omega_{i}}.$ (18)
The interplay between learning rate, batch size, curvature of the loss function, and the variance of fluctuations is then easy to see. We emphasise that the assumption of a single well-defined minimum is not required in the derivation leading to Eq. (16).
We conclude that the stationary distribution after training depends on the ratio of the learning rate and batch size, and not on these quantities separately. This was observed empirically previously Smith-1 ; Smith-2 , but we have demonstrated that it is a direct consequence of stochastic matrix dynamics, when cast in this framework. Ref. DBLP:journals/corr/abs-1710-11029 contains an alternative derivation of the linear scaling rule, using a continuous-time limit of SGD in a weak sense li2017 , while Ref. Sclocchi_2024 contains an application interpreting $\alpha/|{\cal B}|$ as an effective temperature.
The Coulomb gas structure of the stationary distribution has further, universal, consequences. Due to the Coulomb term, eigenvalues repel and the spacing between eigenvalues is non-zero. This effect is captured in the Wigner surmise, which is the distribution for the level spacings $S_{i}=x_{i+1}-x_{i}$, see Eqs. (74, 76) . The spectral density,
$\rho(x)=\left\langle\frac{1}{N}\sum_{i=1}^{N}\delta(x-x_{i})\right\rangle,$ (19)
takes a universal form as well, called the Wigner semicircle, see Eq. (78). In the following we explore both universal and non-universal features in two examples.
III Applications
@@ III.1 Gaussian RBMs
Restricted Boltzmann machines smolensky ; 10.1162/089976602760128018 consist of one visible layer (with $N_{v}$ nodes) and one hidden layer (with $N_{h}$ nodes), see Fig. 1. There are no connections within each layer. The degrees of freedom can be discrete, as in an Ising model, continuous or mixed; Ref. Decelle_2021 is a useful review. We consider Gaussian RBMs, in which both sets of degrees of freedom are quadratic. The ones on the visible layer are collected in an $N_{v}$-dimensional vector $\phi$ and on the hidden layer in an $N_{h}$-dimensional vector $h$. They are coupled bilocally via the $N_{v}\times N_{h}$ weight matrix $W$. We follow the notation of our previous work Aarts:2023uwt .
@@
$E(\phi,h)=\frac{1}{2}\mu^{2}\phi^{T}\phi+\frac{1}{2\sigma_{h}^{2}}(h-\eta)^{T}(h-\eta)-\phi^{T}Wh.$ (21)
@@
$p(\phi)=\int Dh\,p(\phi,h)=\frac{1}{Z}\exp\left(-\frac{1}{2}\phi^{T}K\phi+J^{T}\phi\right),$ (22)
@@
$K=\mu^{2}1\!\!1-\sigma_{h}^{2}WW^{T},\qquad\qquad J=W\eta.$ (23)
Here $1/\mu^{2}$ is the variance on the visible layer, $\sigma_{h}^{2}$ is the variance on the hidden layer, and $\eta$ is a bias, which we will put to zero from now on. This model has been studied in detail from the perspective of lattice field theory (LFT) in Ref. Aarts:2023uwt , which explains the appearance of $\mu^{2}$ as a mass parameter for a scalar field. Its generative power is linked explicitly to the number of hidden nodes $N_{h}$ and the mass parameter $\mu^{2}$, which both act as an ultraviolet regulator, using the terminology familiar from LFT. To match to the general notation of the previous section, we identify $N_{v}=M$ and $N_{h}=N$ and take $N\leq M$.
@@
$\frac{\delta{\cal L}}{\delta W_{ia}}=\sigma_{h}^{2}\sum_{j}\big{(}\langle\phi_{i}\phi_{j}\rangle_{\rm target}-\langle\phi_{i}\phi_{j}\rangle_{\rm model}\big{)}W_{ja}.$ (24)
Here the first two-point function is evaluated using the target data and the second one is the RBM prediction. The dynamics can be analysed semi-analytically Decelle_2021 ; Aarts:2023uwt by performing a singular-value decomposition $W=U\Xi V^{T}$ in which $U$ and $V$ are orthogonal transformations in $M$ and $N$ dimensions respectively, and $\Xi$ is a rectangular matrix with $N$ singular values $\xi_{i}$ ($i=1,\ldots,N$) on the diagonal.
@@
$D_{K}=$
$\mbox{diag}\big{(}\underbrace{\mu^{2}-\sigma_{h}^{2}\xi_{1}^{2},\mu^{2}-\sigma_{h}^{2}\xi_{2}^{2},\ldots,\mu^{2}-\sigma_{h}^{2}\xi_{N}^{2}}_{N},\underbrace{\mu^{2},\ldots,\mu^{2}}_{M-N}\big{)}.$ (26)
@@
$\kappa_{k}=m^{2}+p_{{\rm lat},k}^{2}=m^{2}+2-2\cos\left(\frac{2\pi k}{N_{v}}\right),$ (27)
@@
The numerical training of the RBM is carried out using persistent contrastive divergence (PCD) with mini-batches, see Ref. Aarts:2023uwt for details. The RBM mass parameter is fixed at $\mu^{2}=9$ and $N_{v}=N_{h}=10$. The learning rate and the mini-batch size can be independently varied. For each choice of learning rate and mini-batch size, we have trained 7500 RBMs, to gather statistics. A typical example of the learnt distribution of eigenvalues $x_{i}$ of $X$ is shown in Fig. 3.
@@
III.2 Universal predictions
The task is to describe the distributions shown in Fig. 3 using the concepts from random matrix theory presented above. We start with the universal predictions. Due to the eigenvalue repulsion, degenerate modes cannot be learnt exactly. This can be analysed using the Wigner surmise: for each of the four doubly-degenerate modes we collect data on the level splitting for adjacent eigenvalues, $S_{i}=x_{i+1}-x_{i}$. Here we employed the fact that the non-degenerate eigenvalues are well-separated. The resulting histograms are shown in Fig. 5 (left). These histograms can be fitted with a single-parameter Ansatz – see Eq. (74) – in terms of the mean level splitting $\langle S_{i}\rangle=\sqrt{\pi}\sigma_{i}$. Presenting the same data as a function of $s=S_{i}/\langle S_{i}\rangle$ yields the parameter-free Wigner surmise – see Eq. (76) – leading to a universal description of the level spacing. The linear rise at small $s$ reflects the level repulsion.